OpenStudy (anonymous):

A rocket is launched from atop a 101-foot cliff with an initial velocity of 116 ft/s. a. Substitute the values into the vertical motion formulah=-16t^2+Vot+Ho Let h = 0. b. Use the quadratic formula find out how long the rocket will take to hit the ground after it is launched. Round to the nearest tenth of a second. a.0=-16t^2+101t+116;8s b.0=-16t^2+116t+101;0.8s c.0=-16t^2+101t+116;0.8s d.0=-16t^2+116t+101;8s

OpenStudy (jdoe0001):

$$ \color{blue}{\text{A rocket is launched from atop a 101-foot}}\\ \color{green}{\text{ cliff with an initial velocity of 116 ft/s}}\\ h=-16t^2+\color{green}{V_o}t+\color{blue}{h_o} $$

OpenStudy (jdoe0001):

so, which one do you think is the equation?

OpenStudy (anonymous):

vot

OpenStudy (jdoe0001):

hmm?

OpenStudy (jdoe0001):

well, look at your choices

OpenStudy (anonymous):

so b or d

OpenStudy (jdoe0001):

in the equation $$ \large { \color{green}{V_o} = \text{initial velocity}\\ \color{blue}{h_o} = \text{initial height} } $$

OpenStudy (jdoe0001):

so, what's the initial velocity for the rocket? from what height it initially started traveling?

OpenStudy (anonymous):

116ft/s

OpenStudy (jdoe0001):

and the initial height?

OpenStudy (jdoe0001):

anyhow, you're right, is 101ft, so is either B or D

OpenStudy (jdoe0001):

notice the squared term has a negative number in front of it, that is \(\bf -16t^2+116t+101\ that means the parabola is opening downwards, so it'd look like 

OpenStudy (jdoe0001):

\(\bf -16t^2+116t+101\) that is

OpenStudy (anonymous):

so d would be the appropriate answer

OpenStudy (jdoe0001):

well, we dunno, the rocket will take "t" minutes to hit the ground on the way back down when that happens, that is when the parabola hits the x-axis, y = 0 so \(\bf y = -16t^2+116t+101 \implies 0 =-16t^2+116t+101\)

OpenStudy (jdoe0001):

\(\bf \text{quadratic formula}\\ x= \cfrac{ - b \pm \sqrt { b^2 -4ac}}{2a}\)

OpenStudy (jdoe0001):

now if we plug the values in for that, that'd be a = -16 b = 116 c = 101 \(\bf t= \cfrac{ - 116 \pm \sqrt { 16^2 -4(-16)(101)}}{2(-16)}\)

OpenStudy (jdoe0001):

meh, I have a typo bleh

OpenStudy (jdoe0001):

\(\bf t= \cfrac{ - 16 \pm \sqrt { 16^2 -4(-16)(101)}}{2(-16)}\)

OpenStudy (jdoe0001):

there will be 2 values for "t", one for + root, and one for the - root one number will be negative, you can rule that one out

OpenStudy (jdoe0001):

since it's time and a negative time will mean the rocket was launched today and fell on the ground yesterday, which of course won't make much logical sense hehe

OpenStudy (jdoe0001):

or not yesterday but a few mins ago